# A-TPT: Angular Diversity Calibration Properties for Test-Time Prompt Tuning of Vision-Language Models

### v1

@@ 1 Introduction
To do better prompt calibration, prior works, such as C-TPT (Yoon et al., 2024) and O-TPT (Sharifdeen et al., 2025) have explored methods to encourage dispersion between pairwise textual features, which can be categorized into the following two types: The first type, known as Average Textual Feature Dispersion (ATFD), spreads textual features away from their centroid. However, this can still result in textual features lying closely together (Fig. 2) and cause poor calibration performance (Fig. 2). The second type enforces orthogonality constraints to encourage angular separation, which exploits an auxiliary orthogonal regularization term in the loss function to encourage pairwise textual features as orthogonal as possible. However, we observe that it tends to group textual features closer, particularly when the number of classes $N$ is greater than the embedding dimension $|D|$ (e.g., $\left(N>|D|\right)$, where CLIP’s ViT-B/16 512-d (Radford et al., 2021; Liang et al., 2022) vs. 1000 classes in ImageNet-1k, V2, K) does not guarantee uniformity of angular separation (Fig. 24a). When the number of classes is less than the embedding dimension ($\left(N<|D|\right)$, e.g., classes: 10 in EuroSAT, 37 in OxfordPets, and 47 in DTD), it fails to fully utilize the hyperspherical space of feature points effectively across the hypersphere (Fig. 24b). This eventually leads to poor calibration (Fig. 2). Although these methods increase feature dispersion to some extent, they often overlook the importance of angular diversity. Without sufficient angular separation, prompts may become highly correlated, which limits the model’s ability to generate well-calibrated predictions. Prior work (Wang & Isola, 2020) has shown that uniformity (uniformly distributed feature points on the unit hypersphere) preserves maximal information, closely associated with strong zero-shot CLIP performance.
@@ 2 Related Works
Calibration of deep neural networks. Calibration techniques for deep neural networks can be categorized into two categories: post-hoc and train-time methods. Post-hoc calibration strategies, such as temperature scaling (Guo et al., 2017), platt scaling (Platt et al., 1999), and conformal prediction (Vovk et al., 2005; Lei et al., 2018) calibrate a model’s prediction confidence after training using a held-out validation set. However, these methods rely on access to labeled datasets collected from a distribution similar to the target data (Liu et al., 2022), often impractical in zero-shot and out-of-distribution (OOD) contexts. In contrast, train-time calibration methods integrate a hybrid calibration objective into the training of deep neural networks, with an auxiliary calibration loss as a regularizer in conjunction with the primary training loss. These include techniques (Kumar et al., 2018; Munir et al., 2022; 2023; Yoon et al., 2023) (i.e., for object classification and detection) that incorporate differentiable auxiliary regularization loss functions during training to reduce calibration error for reliable predictions. However, these train-time calibration methods are supervised and require labeled training data, limiting their applicability in the test-time prompt tuning of VLMs without supervision.
@@ 3 Proposed Method
Motivation for angular diversity. Prior works have shown that well-calibrated textual features tend to be more dispersed with L2 distance (Yoon et al., 2024), which is spread farther apart or benefits from greater angular separation with enforced orthogonality constraints (Sharifdeen et al., 2025) in the embedding space. Prior work O-TPT (Sharifdeen et al., 2025) found that lower cosine similarity between textual features corresponds to better model calibration. This suggests that promoting greater angular separation among class-wise text features can improve the calibration of VLMs. However, these analyses overlook a crucial aspect: angular diversity among textual features. We empirically observe that orthogonalization tends to group textual features closer, particularly when the number of classes is greater than the embedding dimension (Fig. 24a), and does not guarantee uniformity of angular separation across all prompt vectors. We hypothesize that promoting angular diversity is more important for better calibration of VLMs than enforcing orthogonality among textual features. Unlike existing techniques that disperse prompts via L2 dispersion (Yoon et al., 2024) or orthogonality constraints (Sharifdeen et al., 2025), our approach explicitly focuses on maximizing the minimum pairwise angular distance among normalized textual features. This numerical method constructs each prompt-induced feature point uniformly distributed on a unit hypersphere. Maximizing the minimum pairwise angular distance between the features ensures each feature vector points in a diverse direction, thereby sharpening class boundaries and improving zero-shot calibration performance in VLMs (Wang & Isola, 2020) during inference.
@@
Comparison of dispersion, orthogonality, and angular diversity. The ATFD (Yoon et al., 2024) objective disperses textual features by maximizing the L2 distance from their centroid, without enforcing pairwise angular separation. In contrast, orthogonality constraints (Sharifdeen et al., 2025) that enforce angular separation by pushing features to be orthogonal to minimize their cosine similarity, tend to group textual features closer together when the number of classes is greater than the embedding dimension $N>|D|$. When the number of classes is lesser $N<|D|$, it fails to fully utilize the hyperspherical space effectively. As a result, neither of these methods guarantees the uniformity of angular separation of features across the hypersphere (Liang et al., 2022). In zero-shot CLIP settings, normalized textual features are constrained to the surface of a unit hypersphere (Wang & Isola, 2020). While ATFD may shift the centroid of the text features toward the center of the hypersphere by adjusting the features, similarly, orthogonality constraints may encourage angular separation by pairwise orthogonalizing the features on the surface. However, these methods may not effectively distribute features uniformly across the hypersphere’s surface. To further validate our findings, we experimented for both where $N<|D|$ (Helber et al., 2018), and $N>|D|$ (Recht et al., 2019) cases. For each data sample, we extract test-time prompt-tuned text features generated by O-TPT and A-TPT (ours), compute pairwise cosine similarities, and plot their mean, as shown in Fig. 4. The results show that O-TPT, which lacks calibration-specific constraints, tends to group textual features closer when $N>|D|$ (O-TPT fails), exhibits lower, but high fluctuations in cosine similarities, reflecting its inconsistent calibration performance. In contrast, our method’s angular diversity consistently produces text features with slightly higher but more consistent cosine similarities, indicating stable, uniform angular separation. Similarly, O-TPT underutilizes hyperspherical space, showing higher, slightly fluctuating cosine similarities when $N<|D|$, while A-TPT shows lower, more consistent cosine similarities, achieving the greatest possible angular separation. That’s why hypersphere offers optimal textual feature separation in A-TPT. (suppl. carries more details)
@@ A.8 Details on the datasets
We evaluate our approach on multiple datasets encompassing both fine-grained classification and natural distribution shift scenarios (suppl. carries details). For fine-grained classification, we conduct experiments using the ImageNet (Deng et al., 2009) and Caltech101 (Fei-Fei et al., 2004) along with diverse datasets across various domains: DTD (Cimpoi et al., 2014) for texture identification, Flower102 (Nilsback & Zisserman, 2008) and OxfordPets (Parkhi et al., 2012) for plants and animal categories, Food101 (Bossard et al., 2014) for food classification, StanfordCars (Krause et al., 2013) and Aircraft (Maji et al., 2013) for transportation classification, SUN397 (Xiao et al., 2010) for scene categorization, UCF101 (Soomro et al., 2012) for human action recognition, and EuroSAT (Helber et al., 2018) for satellite imagery in environmental categorization. For natural distribution shifts, we utilize several ImageNet variants: ImageNet-V2 (Recht et al., 2019) (natural images), ImageNet-A (Hendrycks et al., 2021b) (natural adversarial examples), ImageNet-R (Hendrycks et al., 2021a) (artistic renditions), and ImageNet-Sketch (Wang et al., 2019) (black and white sketches) datasets, as benchmarks for out-of-distribution (OOD) performance evaluation.
@@ A.9 CLIP’s embedding dimension
We employ two CLIP backbones: CLIP ViT-B/16 and CLIP RN50. For all experimental setups, we use test-time prompt tuning (TPT) (Shu et al., 2022) as the primary objective to maximize accuracy while incorporating A-TPT as an auxiliary objective to better the calibration performance as described in Eq. 2. We fix the $\lambda$ as 80.0 for all cases unless otherwise specified. We perform prompt optimization for a single-step update with the AdamW (Loshchilov & Hutter, 2017) optimizer and set the learning rate to 5e-3. We initialize the prompt embeddings with hard prompts, following C-TPT (Yoon et al., 2024) and all other settings following the standard TPT (Shu et al., 2022) configurations. We conduct all experiments with a batch size of 64 on a single NVIDIA Quadro RTX 6000 GPU (24GB memory).
@@ A.16 Why hypersphere offers optimal textual feature separation in A-TPT?
Firstly, based on the Tammes problem (Tammes, 1930), the uniform distribution (best-packing) means the distance between two points is maximized; therefore, while A-TPT builds upon the TPT framework, it introduces a novel angular diversity that is boosted to the utmost extent. Angular diversity (AD) focuses on the uniform distribution — maximizing the angular distance between textual features corresponding to their actual class labels, thereby maximizing the inter-class feature separability for better prompt calibration. Prior work (Wang & Isola, 2020) has shown that uniformity (uniformly distributed feature points on the unit hypersphere) preserves maximal information, closely associated with strong zero-shot CLIP performance. Therefore, we argue that uniformly distributed textual features are the way to ensure better class separation and improve VLM calibration. Unlike orthogonality constraints pursue orthogonality for all the pairwise textual features and suffer from poor calibration when $N>|D|$, our numerical optimization (A-TPT) is robust with negligible computational overhead across both $N>|D|$ and $N<|D|$ settings. Secondly, inspired by insights from ArcFace (Deng et al., 2019), Figs. 24, and Appendix A.16 of the main paper demonstrate that AD gets the greatest minimum pairwise angular distance across all $N>|D|$ & $N<|D|$ cases, and therefore the most diverse prompt vectors. Thirdly, as discussed in the paper, orthogonal regularization tends to group textual features closer, especially when the number of classes is greater than the embedding dimension. Fourthly, we argue that the improvement over other methods is significant.

### v2

@@ 1 Introduction
To do better prompt calibration, prior works, such as C-TPT (Yoon et al., 2024) and O-TPT (Sharifdeen et al., 2025) have explored methods to encourage dispersion between pairwise textual features, which can be categorized into the following two types: The first type, known as Average Textual Feature Dispersion (ATFD), spreads textual features away from their centroid. However, this can still result in textual features lying closely together (Fig. 2) and cause poor calibration performance (Fig. 2). The second type enforces orthogonality constraints to encourage angular separation, which exploits an auxiliary orthogonal regularization term in the loss function to encourage pairwise textual features as orthogonal as possible. However, we observe that it tends to group textual features closer, particularly when the number of classes $N$ is greater than the embedding dimension $|D|$ (e.g., $\left(N>|D|\right)$, where CLIP’s ViT-B/16 512-d (Radford et al., 2021; Liang et al., 2022) vs. 1000 classes in ImageNet-1k, V2, K) does not guarantee uniformity of angular separation (Fig. 24a). When the number of classes is less than the embedding dimension ($\left(N<|D|\right)$, e.g., classes: 10 in EuroSAT, 37 in OxfordPets, and 47 in DTD), it fails to fully utilize the hyperspherical space of feature points effectively across the hypersphere (Fig. 24b). This eventually leads to poor calibration (Fig. 2). Although these methods increase feature dispersion to some extent, they often overlook the importance of angular diversity. Without sufficient angular separation, prompts may become highly correlated, which limits the model’s ability to generate well-calibrated predictions. Prior work (Wang and Isola, 2020) has shown that uniformity (uniformly distributed feature points on the unit hypersphere) preserves maximal information, closely associated with strong zero-shot CLIP performance.
@@ 2 Related Works
Calibration of deep neural networks. Calibration techniques for deep neural networks can be categorized into two categories: post-hoc and train-time methods. Post-hoc calibration strategies, such as temperature scaling (Guo et al., 2017), platt scaling (Platt and others, 1999), and conformal prediction (Vovk et al., 2005; Lei et al., 2018) calibrate a model’s prediction confidence after training using a held-out validation set. However, these methods rely on access to labeled datasets collected from a distribution similar to the target data (Liu et al., 2022), often impractical in zero-shot and out-of-distribution (OOD) contexts. In contrast, train-time calibration methods integrate a hybrid calibration objective into the training of deep neural networks, with an auxiliary calibration loss as a regularizer in conjunction with the primary training loss. These include techniques (Kumar et al., 2018; Munir et al., 2022; 2023; Yoon et al., 2023) (i.e., for object classification and detection) that incorporate differentiable auxiliary regularization loss functions during training to reduce calibration error for reliable predictions. However, these train-time calibration methods are supervised and require labeled training data, limiting their applicability in the test-time prompt tuning of VLMs without supervision.
@@ 3 Proposed Method
Motivation for angular diversity. Prior works have shown that well-calibrated textual features tend to be more dispersed with L2 distance (Yoon et al., 2024), which is spread farther apart or benefits from greater angular separation with enforced orthogonality constraints (Sharifdeen et al., 2025) in the embedding space. Prior work O-TPT (Sharifdeen et al., 2025) found that lower cosine similarity between textual features corresponds to better model calibration. This suggests that promoting greater angular separation among class-wise text features can improve the calibration of VLMs. However, these analyses overlook a crucial aspect: angular diversity among textual features. We empirically observe that orthogonalization tends to group textual features closer, particularly when the number of classes is greater than the embedding dimension (Fig. 24a), and does not guarantee uniformity of angular separation across all prompt vectors. We hypothesize that promoting angular diversity is more important for better calibration of VLMs than enforcing orthogonality among textual features. Unlike existing techniques that disperse prompts via L2 dispersion (Yoon et al., 2024) or orthogonality constraints (Sharifdeen et al., 2025), our approach explicitly focuses on maximizing the minimum pairwise angular distance among normalized textual features. This numerical method constructs each prompt-induced feature point uniformly distributed on a unit hypersphere. Maximizing the minimum pairwise angular distance between the features ensures each feature vector points in a diverse direction, thereby sharpening class boundaries and improving zero-shot calibration performance in VLMs (Wang and Isola, 2020) during inference.
@@
Comparison of dispersion, orthogonality, and angular diversity. The ATFD (Yoon et al., 2024) objective disperses textual features by maximizing the L2 distance from their centroid, without enforcing pairwise angular separation. In contrast, orthogonality constraints (Sharifdeen et al., 2025) that enforce angular separation by pushing features to be orthogonal to minimize their cosine similarity, tend to group textual features closer together when the number of classes is greater than the embedding dimension $N>|D|$. When the number of classes is lesser $N<|D|$, it fails to fully utilize the hyperspherical space effectively. As a result, neither of these methods guarantees the uniformity of angular separation of features across the hypersphere (Liang et al., 2022). In zero-shot CLIP settings, normalized textual features are constrained to the surface of a unit hypersphere (Wang and Isola, 2020). While ATFD may shift the centroid of the text features toward the center of the hypersphere by adjusting the features, similarly, orthogonality constraints may encourage angular separation by pairwise orthogonalizing the features on the surface. However, these methods may not effectively distribute features uniformly across the hypersphere’s surface. To further validate our findings, we experimented for both where $N<|D|$ (Helber et al., 2018), and $N>|D|$ (Recht et al., 2019) cases. For each data sample, we extract test-time prompt-tuned text features generated by O-TPT and A-TPT (ours), compute pairwise cosine similarities, and plot their mean, as shown in Fig. 4. The results show that O-TPT, which lacks calibration-specific constraints, tends to group textual features closer when $N>|D|$ (O-TPT fails), exhibits lower, but high fluctuations in cosine similarities, reflecting its inconsistent calibration performance. In contrast, our method’s angular diversity consistently produces text features with slightly higher but more consistent cosine similarities, indicating stable, uniform angular separation. Similarly, O-TPT underutilizes hyperspherical space, showing higher, slightly fluctuating cosine similarities when $N<|D|$, while A-TPT shows lower, more consistent cosine similarities, achieving the greatest possible angular separation. That’s why hypersphere offers optimal textual feature separation in A-TPT. (suppl. carries more details)
@@ A.8 Details on the datasets
We evaluate our approach on multiple datasets encompassing both fine-grained classification and natural distribution shift scenarios (suppl. carries details). For fine-grained classification, we conduct experiments using the ImageNet (Deng et al., 2009) and Caltech101 (Fei-Fei et al., 2004) along with diverse datasets across various domains: DTD (Cimpoi et al., 2014) for texture identification, Flower102 (Nilsback and Zisserman, 2008) and OxfordPets (Parkhi et al., 2012) for plants and animal categories, Food101 (Bossard et al., 2014) for food classification, StanfordCars (Krause et al., 2013) and Aircraft (Maji et al., 2013) for transportation classification, SUN397 (Xiao et al., 2010) for scene categorization, UCF101 (Soomro et al., 2012) for human action recognition, and EuroSAT (Helber et al., 2018) for satellite imagery in environmental categorization. For natural distribution shifts, we utilize several ImageNet variants: ImageNet-V2 (Recht et al., 2019) (natural images), ImageNet-A (Hendrycks et al., 2021b) (natural adversarial examples), ImageNet-R (Hendrycks et al., 2021a) (artistic renditions), and ImageNet-Sketch (Wang et al., 2019) (black and white sketches) datasets, as benchmarks for out-of-distribution (OOD) performance evaluation.
@@ A.9 CLIP’s embedding dimension
We employ two CLIP backbones: CLIP ViT-B/16 and CLIP RN50. For all experimental setups, we use test-time prompt tuning (TPT) (Shu et al., 2022) as the primary objective to maximize accuracy while incorporating A-TPT as an auxiliary objective to better the calibration performance as described in Eq. 2. We fix the $\lambda$ as 80.0 for all cases unless otherwise specified. We perform prompt optimization for a single-step update with the AdamW (Loshchilov and Hutter, 2017) optimizer and set the learning rate to 5e-3. We initialize the prompt embeddings with hard prompts, following C-TPT (Yoon et al., 2024) and all other settings following the standard TPT (Shu et al., 2022) configurations. We conduct all experiments with a batch size of 64 on a single NVIDIA Quadro RTX 6000 GPU (24GB memory).
@@ A.16 Why hypersphere offers optimal textual feature separation in A-TPT?
Firstly, based on the Tammes problem (Tammes, 1930), the uniform distribution (best-packing) means the distance between two points is maximized; therefore, while A-TPT builds upon the TPT framework, it introduces a novel angular diversity that is boosted to the utmost extent. Angular diversity (AD) focuses on the uniform distribution — maximizing the angular distance between textual features corresponding to their actual class labels, thereby maximizing the inter-class feature separability for better prompt calibration. Prior work (Wang and Isola, 2020) has shown that uniformity (uniformly distributed feature points on the unit hypersphere) preserves maximal information, closely associated with strong zero-shot CLIP performance. Therefore, we argue that uniformly distributed textual features are the way to ensure better class separation and improve VLM calibration. Unlike orthogonality constraints pursue orthogonality for all the pairwise textual features and suffer from poor calibration when $N>|D|$, our numerical optimization (A-TPT) is robust with negligible computational overhead across both $N>|D|$ and $N<|D|$ settings. Secondly, inspired by insights from ArcFace (Deng et al., 2019), Figs. 24, and Appendix A.16 of the main paper demonstrate that AD gets the greatest minimum pairwise angular distance across all $N>|D|$ & $N<|D|$ cases, and therefore the most diverse prompt vectors. Thirdly, as discussed in the paper, orthogonal regularization tends to group textual features closer, especially when the number of classes is greater than the embedding dimension. Fourthly, we argue that the improvement over other methods is significant.